OpenStudy (anonymous):

Find the zeros of the function by using the Quadratic Formula. f(x) = x2 + 3x - 7

OpenStudy (anonymous):

Please explain! I thought it would be a simple answer but my choices seem complex..So confused.

OpenStudy (anonymous):

the quadratic formula is \(\large{x={-b\pm\sqrt{b^2-4ac}\over2a}}\) where \(f(x)=ax^2+bx+c\)

OpenStudy (anonymous):

so for your equation \(f(x)=\color{red}{1}x^2\color{blue}{+3}x\color{green}{-7}\) the quadratic is to be solved as such:\[x={-\color{red}{3}\pm\sqrt{\color{red}{3}^2-4(\color{blue}{1})(\color{green}{-7})}\over2(\color{blue}{1})}\]now just solve through, and be mindful of the "\(\pm\)" sign, which means plus or minus, so you'll end up with two answers :)